OpenStudy (anonymous):

Use Gausses approach to find the following sums a. 1+2+3+4+...+102

OpenStudy (zepp):

Add the first term, and the last term, and multiply by, the sequence median. :D

OpenStudy (anonymous):

Gauss approach is add forward and backward to get twice the sum . Then divide by two

OpenStudy (zepp):

The median is 51, I'll just go \[\large \sum_{x=1}^{102}x=51(1+102)=5253\]

OpenStudy (anonymous):

so 1 +102=103, 2 +101 =1-3, 3+100=103....how long do i use it?

OpenStudy (zepp):

51 times!

OpenStudy (anonymous):

Hey Zep can you explain how you got 51?

OpenStudy (anonymous):

you dont have to be rude, i had already sent my post before I saw your response

OpenStudy (zepp):

I wasn't rude haha :P

OpenStudy (anonymous):

I just want to understand how you got 51 so quickly so i can practice this method :)

OpenStudy (zepp):

As I said earlier, you can simply add up the first term to the second term and multiply this sum by the median of this sequence, and voila! (Median of the sequence could be found be simply take the last number of the sequence, and divide it by 2)

OpenStudy (zepp):

In this case, you do 102/2 = 51 :)

OpenStudy (anonymous):

you mean the first term by the last term and divide by two? so 1+3+5+7+...1001? the sum =502002??

OpenStudy (zepp):

Um, that would be 251 001

OpenStudy (anonymous):

(1001+1)/2=501?

OpenStudy (anonymous):

501(1+1001)=

OpenStudy (zepp):

Well, these are odd numbers, we can write them as 2n-1 \[\large \sum_{x=1}^{501}2n-1\]

OpenStudy (anonymous):

ohhhhhh lol...thank you! i'll keep practicing..

OpenStudy (zepp):

Which is what you got above, divided by 2

OpenStudy (zepp):

"you mean the first term by the last term and divide by two? so 1+3+5+7+...1001? the sum =502002??" 502 002/2 = 251 001 :)

OpenStudy (anonymous):

ohh i didnt finish the problem lol